OpenStudy (anonymous):

Find the slope y^2-y^1 m = -------- x^2-x^1 of a line passing through the given points. (-4,2) and (1,-3)

OpenStudy (anonymous):

Slope is 1.

OpenStudy (anonymous):

I thought so

OpenStudy (waleed_imtiaz):

Slope is -1 I think so... Because -3-2/1+4=-5/5=-1

OpenStudy (anonymous):

waleed is right. I missed a negative

OpenStudy (waleed_imtiaz):

hm... Ya.. no problem....